OpenStudy (ebonyserrano):

Chau's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Chau $5.70 per pound, and type B coffee costs $4.35 per pound. This month, Chau made 179 pounds of the blend, for a total cost of $886.65. How many pounds of type B coffee did he use?

OpenStudy (ebonyserrano):

@Will.H

OpenStudy (revlogic):

for coffee A = 5.70/lb for coffee B = 4.35/lb x = # of lb for coffee A y = # of lb for coffee B x(5.70) + y(4.35) = 886.65 We also know that combine pounds for both coffee A and B is 179 x + y = 179 -------- Now just solve for x for the first equation and substitute that value to 2nd equation. Afterward, just solve for y and that's the value for how many pounds for coffeeB.

OpenStudy (ebonyserrano):

i got 120

OpenStudy (revlogic):

Let me check: x(5.70) + y(4.35) = 886.65 5.70x = 886.65 -4.35y x = (886.65-4.35)/5.70 now substitute that x value into equation #2: (886.65-4.35y)/5.70 + y = 179 http://www.wolframalpha.com/input/?i=(886.65+-4.35y)%2F5.70++%2B+y+%3D+179 y = 99

OpenStudy (revlogic):

You can double check the work: 99 * 4.35 = 430.65 88 * 5.7 =456 so, 430.65 + 456 ===> 886.65

OpenStudy (revlogic):

Edit, 88 is actually 80

OpenStudy (ebonyserrano):

so it would be 80 pounds?

OpenStudy (revlogic):

No, you want value for coffee B, so it's value for y y= 99

OpenStudy (ebonyserrano):

@Will.H would it be 66 pounds then?

OpenStudy (dumbcow):

@Ebonyserrano , refer to previous posts You want type B which is represented by "y" ... y = 99